OpenStudy (anonymous):

56. Using Newton’s second law of motion, if the force on an object is 20 N, and its mass is 5 g, what is its acceleration? (3 points)

OpenStudy (anonymous):

Mathematical formulation of Newton's second law: F = m * a. I think mass should be in kgs. So answer should be a = F/m = 20/0.05 = 2000/5 = 400 m/s^2

OpenStudy (shane_b):

Almost right :) \[F=ma=>a=\frac{F}{m}=\frac{20N}{0.005kg}=4,000~m/s^2\]